Can you help with the following problems and thank you for your help.
1. -10 + (-3 - 5) =
2. 5 - (-8 + 7) =
3. (3 + -4) + 2 =
4. 9 + (3 - 7) =
5. 1 - (6 + 4)=

Answer by brandonpark2889(31)

Can you help with the following problems and thank you for your help.
1. -10 + (-3 - 5) =
2. 5 - (-8 + 7) =
3. (3 + -4) + 2 =
4. 9 + (3 - 7) =
5. 1 - (6 + 4)=

These answers require the order of operation. The 1st one:
1. -10 + (-3 - 5). You first add whatever it is in the parenthesis, which is -3 and -5. You can easily combine them by adding these, equaling -8. Lastly, -10 + -8 is -18.
2. 5 - (-8 + 7). Whatever in the parenthesis first. You will combine -8 and 7, which equals -1. Now it will be left over with 5 - -1. Make these subtraction signs positive, which equals 6.
3. (3 + -4) + 2. First, the parenthesis: 3 and -4. They equal to -1, and add it with 2 to make 1.
4. 9 + (3 - 7). Parenthesis first: 3 - 7 = -4, and combine it with 9 to make 5.
5. 1 - (6 + 4). This is little trickier. Combine 6 with 4. Then it will be 1 - 10, meaning the added numbers are actually negative. 1 - 10 would equal -9.
[1] -10 + (-3 - 5) = x
-10 + -8 = x
-18 = x
[2] 5 - (-8 + 7) = x
5 - - 1 = x
5 + 1 = x
6 = x
[3] (3 + -4) + 2 = x
-1 + 2 = x
1 = x
[4] 9 + (3 - 7) = x
9 + -4 = x
5 = x
[5] 1 - (6 + 4) = x
1 - 10 = x
-9 = x
Alright? Good luck.